4
(4) Evaluate S 3x²(2x°-3) dx.
Answer #1

substitute u = 2.13 - 3 then do = 2(3x?) + de a 1151 dr = 1 1 46 3:2 (2:23 -3)5 d.x = = 5+1 5+1) + 5 +1 += = 7 2 + substitute